OpenStudy (anonymous):

please help solve: integral sinx^(5)dx

OpenStudy (anonymous):

let \[u=\cos(x)\] so \[du=-\sin(x) dx\] thus \[\int\limits_{}^{} \sin^5 (x) dx = \int\limits_{}^{} \sin^4 (x) \sin(x) dx= \int\limits_{}^{} (1- \cos^2(x))^2 \sin(x) dx\] =\[-\int\limits_{}^{}(1-u^2)^2 du = -\int\limits_{}^{} (1-2u^2+u^4) du\] I guess you can do the last integral, just remeber to replace u by cos(x) after you done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u r welcome